Can two or more three-dimensional solid figures have the same volume? Why or why not?


1
Expert's answer
2022-04-21T09:56:14-0400

Yes, they can.

For example, cone and cylinder. Take a cylinder with a base radius of 1 meter and a height of 1 meter

"V = \pi r^2h"

"V = \pi"

And take a cone with a base radius of 1 meter and a height of 3 meters

"V = {1 \over 3}\pi r^2h = {1 \over 3}*3\pi=\pi"

"V = \pi"

Both of these figures have the same volume.